# DefenderBench: A Toolkit for Evaluating Language Agents in Cybersecurity Environments

### v2

@@ 1 Introduction
LLMs (Touvron et al., 2023a, b; OpenAI, 2023) have demonstrated impressive capacities for understanding and generating natural language. To better leverage LLMs for real-world problem-solving, recent works (Zhao et al., 2024; Park et al., 2023; Wang et al., 2023; Wu et al., 2024a) have integrated LLMs into agentic frameworks, enabling them to perform tasks by interacting with an environment (ecosystem), communicating with multiple agents, and breaking down complex tasks into simpler ones to achieve a higher degree of automation. Recent studies have shown that LLM-based agentic systems effectively handle diverse tasks such as software development (Qian et al., 2024), document-level machine translation (Wu et al., 2024b), and fact-checking (Du et al., 2024). Several concurrent studies have introduced evaluation benchmarks to better assess the capabilities of LLM-based agentic systems, including AgentBench (Liu et al., 2024a) for system and database operations, MLAgentBench (Huang et al., 2024) for machine learning research, SWE-bench (Jimenez et al., 2024) for software development, SmartPlay (Wu et al., 2024c) for games, and WebArena (Zhou et al., 2024) for web workflows. However, how LLM-based agents address cybersecurity-related tasks remains underexplored. Although some contemporaneous works have begun developing evaluation benchmarks for LLM agents in cybersecurity, such as Cybench (Zhang et al., 2024a) for Capture The Flag challenges, CyberMetric (Tihanyi et al., 2024) for cybersecurity knowledge question answering, and CyberSecEval (Bhatt et al., 2024) for code vulnerability detection and exploitation, they focus solely on one or a few specific cybersecurity tasks.
To further explore the capabilities of LLM agents in cybersecurity and enhance fairness of model comparisons and reproducibility, we introduce DefenderBench, a toolkit for evaluating LLM-based agents on cybersecurity tasks. As a dual-use technology (Zhang et al., 2024a; Biden, 2023), LLM agents for cybersecurity are evaluated on three types of tasks: offense, defense, and cybersecurity knowledge understanding. For offense tasks, we implement a text-based wrapper around a network intrusion environment with various configurations. For defense tasks, we include malicious content detection, code vulnerability detection, and code vulnerability fixing. Additionally, we incorporate a multiple-choice question-answering task to assess LLM agents’ understanding of cybersecurity knowledge. Inspired by existing LLM agentic frameworks (Wu et al., 2024c; Liu et al., 2024a; Wei et al., 2022), we introduce an agent baseline to benchmark different LLMs on these cybersecurity tasks. We evaluate several LLMs including open-weight models from the Llama (Dubey et al., 2024) and Phi (Abdin et al., 2024) families, along with proprietary models such as the GPTs (OpenAI, 2023) and Claudes [^1] https://www.anthropic.com/claude. Our experiments show that Claude-3.7-sonnet is the best-performing LLM with a DefenderBench score of 81.65.
@@ 2 Related Work
LLM for Cybersecurity. With our growing reliance on digital and interconnected systems and the increasing sophistication of cyber threats (Thakur et al., 2015), cybersecurity has become a critical area of focus. Cybersecurity encompasses a comprehensive range of practices, tools, and strategies aimed at protecting computer systems, networks, and data from unauthorized access, attacks, damage, or disruptions (Li and Liu, 2021; Zhang et al., 2024c). Traditional cybersecurity approaches, such as rule-based systems, struggle to keep pace with rapidly evolving cyber threats. With advancements in LLMs, efforts have been made to leverage LLMs to address cybersecurity challenges. For instance, domain-specific datasets have been curated to fine-tune LLMs for tasks such as program repair (Silva et al., 2023), cybersecurity training (Zhang et al., 2023), network security (Rigaki et al., 2024) and secure code generation (Mechri et al., 2025). Additionally, LLM agents have been employed in tasks like website hacking (Fang et al., 2024b), code vulnerability exploitation (Fang et al., 2024a), debugging (Lee et al., 2024), and penetration testing (Deng et al., 2023). In this paper, we focus on developing a standardized toolkits for evaluating LLM agents.
@@
LLM Agent Benchmark. To evaluate the capabilities of LLM agents, several benchmarks have been developed. AgentBench (Liu et al., 2024a) assesses LLMs across five diverse environments, including operating systems and databases, to evaluate reasoning and decision-making abilities. MLAgentBench (Huang et al., 2024) focuses on machine learning experimentation tasks, testing agents on tasks ranging from improving model performance to addressing research problems. SWE-bench (Jimenez et al., 2024) evaluates LLMs on real-world software issues sourced from GitHub, requiring models to generate patches that resolve described problems. SmartPlay (Wu et al., 2024c) introduces a suite of games to test various capabilities of LLMs, such as planning and spatial reasoning. WebArena (Zhou et al., 2024) provides a realistic web environment for building autonomous agents, enabling the assessment of LLMs in web-based tasks.
@@
Cybersecurity-Specific Benchmarks. In the cybersecurity domain, specialized benchmarks have been introduced. Cybench (Zhang et al., 2024a) offers a framework for evaluating LLM agents on 40 professional-level Capture The Flag (CTF) tasks, encompassing a range of difficulties and scenarios. CyberMetric (Tihanyi et al., 2024) presents a benchmark dataset based on retrieval-augmented generation to assess LLMs’ cybersecurity knowledge. SecEval (Li et al., 2023) provides over 2,000 multiple-choice questions across various cybersecurity domains to evaluate foundation models’ knowledge. CyberSecEval (Bhatt et al., 2024) focuses on code vulnerability detection and exploitation, offering a comprehensive suite for assessing LLMs in secure coding tasks. These benchmarks facilitate targeted evaluations of LLMs in cybersecurity contexts. The closest work to ours is CyberBench Liu et al. (2024b), a benchmark focusing on Natural Language Processing (NLP) tasks related to cybersecurity.
@@
DefenderBench. We introduce DefenderBench, a toolkit designed to evaluate LLM agents in interactive cybersecurity environments. Unlike existing benchmarks mentioned above that focus on specific tasks or domains, DefenderBench encompasses a broad range of cybersecurity-related tasks, covering offense, defense, and knowledge understanding. By integrating insights from general agent benchmarks and adversarial evaluation frameworks, DefenderBench aims to provide a comprehensive assessment platform for LLMs in cybersecurity contexts.
@@ 3.1 Computer Network Intrusion Simulation
In order to protect computer networks against attacks, many organizations conduct red-team network intrustion to proactively detect and remediate vulnerabilites before attackers do. We leverage the network intrusion simulation tool CyberBattleSim (CBS) (Team., 2021) to evaluate the ability of LLM agents to identify vulnerabilities in a network. CyberBattleSim is parameterized by a fixed topology and a set of node vulnerabilities that agents can exploit to move laterally within the network. The goal of the attacker is to take ownership of the network by exploiting vulnerabilities in the computer nodes. We convert CyberBattleSim into a text-based game (Côté et al., 2019) which describes the currently discovered network as some structured text (i.e., JSON) and provides textual feedback in response to the agent’s actions. There are three action types for an attacker to interact with the network:
@@ 3.2 Malicious Content Detection
Malicious-Text: for this task, we utilize the dataset processed by Alvarado (2024). [^2] https://huggingface.co/datasets/ealvaradob/phishing-dataset This dataset incorporates two data sources, namely email and text messages, for malicious content detection. The entire dataset contains 20,137 samples labeled as $\{malicious,legitimate\}$. We follow the split of Alvarado (2024), using 80% of the data as the training set and 20% as the test set. To reduce the cost of performing LLMs on our benchmark, we further randomly select 500 samples from the test split as our official test set in the benchmark. Additionally, we select 10 samples per class as the few-shot sampling pool for in-context learning (ICL) Brown et al. (2020). The metric used is the macro-F1 score.
Malicious-Web: This task assesses the ability of LLM agents to discriminate phishing from benign web sites. We use the Phishing Websites Dataset (Ariyadasa et al., 2021) as preprocessed by Alvarado (2024) for malicious website detection. We also discard 144 samples which contain less than 100 characters as they are mostly outliers (e.g. page failed to load). The resulting dataset (15,612 samples) includes 10,220 labeled as legitimate and 5,392 as malicious. We follow the same 80%-20% split as Alvarado (2024) and further uniformly subsample 500 test samples as our test set and 10 training samples per class as the few-shot sampling pool. We report the macro-F1 score for this task.
@@ 3.3 Cyber Threat Intelligence (CTI)
MCQA: This task assesses the ability of an LLM agent to understand recent threat intelligence and apply it to challenging questions. A multiple-choice question answering task that uses the CTI-MCQA dataset introduced by Alam et al. (2024). This dataset originally contains 2,500 questions, each associated with a CTI-related webpage or document. After filtering out questions linked to inaccessible webpage or document, we obtained 2,338 samples. We then randomly downsample and split these into a test set (500 questions) and a few-shot sampling pool (20 samples). Each question has four options, with only one correct answer. The metric for this task is macro-F1.
@@ 3.4 Code Vulnerability Detection
Vulnerable-CG: This task assesses the ability of LLM agents to detect vulnerabilities in code. We use the code vulnerability detection dataset included in CodeXGLUE (Lu et al., 2021), which is split into training (21,854 samples), validation (2,732 samples), and test sets (2,732 samples). Each sample is a C language function annotated with the label ‘vulnerable’ or ‘non-vulnerable’. Our test samples are 500 randomly selected samples from their test set. We also provide 10 training samples per class as the few-shot sampling pool. The agent’s performance is reported using the macro-F1 score.
Vulnerable-DV: we also include the Devign (Zhou et al., 2019) dataset for code vulnerability detection in our benchmark. Zhou et al. (2019) released two projects, FFmpeg and Qemu, comprising a total of 27,318 samples. We randomly sample 500 samples for our test set. Similarly, we include 10 training samples per class as the few-shot sampling pool and report the macro-F1 score as the evaluation metric.
@@ 3.5 Code Vulnerability Fixing
CVEFix: we use the CVEFix dataset (Bhandari et al., 2021) for the vulnerability fixing task. The original dataset contains 12,107 vulnerability fixing commits across 4,249 open-source projects. The dataset includes the source code before and after the changes. We only extract commits with the following conditions: (a) single method modification; (b) the commit is associated to a single CVE (Common Vulnerabilities and Exposures); (c) the programming languages is either: C, C++, Go, Java, JavaScript, PHP, Python, or Rust. As a result, we obtained 240 samples. We use all the samples as the test set for our benchmark. For this task, we provide the method’s source code before the commit and ask the agent to generate a new method that fixes any vulnerability. We report the CodeBLEU score (Ren et al., 2020) between the generated method and the method after the commit.
@@ 4.1 Modules
Data Preprocessing. The DefenderBench toolkit automatically downloads the required datasets from their respective sources, shuffles the samples randomly according to a fixed random seed, and splits them into a test set and a few-shot sample pool for in-context learning. Once preprocessed, the datasets are cached locally. For network intrusion simulation, we install CyberBattleSim (Team., 2021) as a dependency.
@@
Task Environment. For each task, we set up a task environment that provides task-specific instructions (shown in Table 1), defines the action space for the agent, loads the relevant datasets and constructs few-shot examples if few-shot in-context learning is being conducted (more on this in section 5.3). For the detection, MCQA, and code-fixing tasks, each episode involves presenting the agent with a test sample. Each episode can run for up to five steps. If the agent fails to respond with the expected format, a feedback message is provided and the agent can try again until the episode ends. For the network intrusion task, each episode begins with an initialized network and can run for up to 100 steps to compromise the entire network. The LLM agent interacts with the task environment by providing a text action and the environment provides an observation in return. The observation describes the result of the given action and indicates whether the task has been completed. Additionally, the environment maintains a history of the actions taken by the agent and the corresponding feedback. The history can be provided to the agent as part of its context.
@@
Agent Interface. Our DefenderBench is equipped with an LLM agent interface that enables users to integrate both open- and closed-weight LLMs. Users can also seamlessly incorporate their own agentic system to perform the tasks.
@@
Execution. To evaluate LLM agents on DefenderBench, users can install our toolkit as a Python library. Through a terminal command, users can run all tasks or specify a particular task by using its shorthand name. Additionally, users can choose which LLM to use for the baseline agent. We have also integrated the Weights and Biases library into DefenderBench, [^3] https://wandb.ai/ enabling users to track and visualize their results seamlessly.
@@
Metrics. We report on each task using its original metric as described in Section 3. Inspired by previous evaluation benchmarks like GLUE (Wang et al., 2019), we define a global metric called DefenderBench score, which represents the unweighted average of all task-specific metrics. The DefenderBench score provides an overall indication of performance on cybersecurity tasks.
@@
Baseline Agent. To evaluate the out-of-the-box capability of LLMs in solving cybersecurity tasks, we experiment with a baseline agent with minimal scaffolding in this paper. As illustrated in Figure 1, we begin by providing to the agent a task instruction that explains the task, specifies the response format, and defines the action space. Table 1 shows the task instructions. At each step, the agent is given the trajectory of its prior actions along with the corresponding observations from the environment. At each step, the agent is asked to produce an action in the required format, which is then sent to the task environment to obtain an action observation. Based on this observation, we determine whether the episode should be terminated. If the episode continues, the observation is added to the system prompt as part of the historical trajectory.
@@ 5.1 Backbone LLMs
In our experiments, we use a variety of LLMs as the backbone of our agent. These include (1) open-weight models (Llama 3.1 (Dubey et al., 2024), Llama 3.2, Llama 3.3, and Phi-3.5 (Abdin et al., 2024)), (2) proprietary models (GPT-3.5, GPT-4-turbo, GPT-4o, GPT-4o-mini, Claude-3.5-haiku, and Claude-3.5-sonnet, Claude-3.7-sonnet), and (3) proprietary reasoning models (o1, o1-mini, o3, o4-mini, GPT-4.1, GPT-4.1-mini, and GPT-4.1-nano, Claude-3.7-sonnet-think).

### v4

@@ 1 Introduction
LLMs [1, 2, 3] have demonstrated impressive capacities for understanding and generating natural language. To better leverage LLMs for real-world problem-solving, recent works [4, 5, 6, 7] have integrated LLMs into agentic frameworks, enabling them to perform tasks by interacting with an environment (ecosystem), communicating with multiple agents, and breaking down complex tasks into simpler ones to achieve a higher degree of automation. Recent studies have shown that LLM-based agentic systems effectively handle diverse tasks such as software development [8], document-level machine translation [9], and fact-checking [10]. Several concurrent studies have introduced evaluation benchmarks to better assess the capabilities of LLM-based agentic systems, including AgentBench [11] for system and database operations, MLAgentBench [12] for machine learning research, SWE-bench [13] for software development, SmartPlay [14] for games, and WebArena [15] for web workflows. However, how LLM-based agents address cybersecurity-related tasks remains underexplored. Although some contemporaneous works have begun developing evaluation benchmarks for LLM agents in cybersecurity, such as Cybench [16] for Capture The Flag challenges, CyberMetric [17] for cybersecurity knowledge question answering, and CyberSecEval [18] for code vulnerability detection and exploitation, they focus solely on one or a few specific cybersecurity tasks.
To further explore the capabilities of LLM agents in cybersecurity and enhance fairness of model comparisons and reproducibility, we introduce DefenderBench, a toolkit for evaluating LLM-based agents on cybersecurity tasks. As a dual-use technology [16, 19], LLM agents for cybersecurity are evaluated on three types of tasks: offense, defense, and cybersecurity knowledge understanding. For offense tasks, we implement a text-based wrapper around a network intrusion environment with various configurations. For defense tasks, we include malicious content detection, code vulnerability detection, and code vulnerability fixing. Additionally, we incorporate a multiple-choice question-answering task to assess LLM agents’ understanding of cybersecurity knowledge. Inspired by existing LLM agentic frameworks [14, 11, 20], we introduce an agent baseline to benchmark different LLMs on these cybersecurity tasks. We evaluate several LLMs including open-weight models from the Llama [21] and Phi [22] families, along with proprietary models such as the GPTs [3] and Claudes [^1] https://www.anthropic.com/claude. Our experiments show that Claude-3.7-sonnet is the best-performing LLM with a DefenderBench score of 81.65.
@@ 2 Related Work
With our growing reliance on digital and interconnected systems and the increasing sophistication of cyber threats [23], cybersecurity has become a critical area of focus. Cybersecurity encompasses a comprehensive range of practices, tools, and strategies aimed at protecting computer systems, networks, and data from unauthorized access, attacks, damage, or disruptions [24, 25]. Traditional cybersecurity approaches, such as rule-based systems, struggle to keep pace with rapidly evolving cyber threats. With advancements in LLMs, efforts have been made to leverage LLMs to address cybersecurity challenges. For instance, domain-specific datasets have been curated to fine-tune LLMs for tasks such as program repair [26], cybersecurity training [27], network security [28] and secure code generation [29]. Additionally, LLM agents have been employed in tasks like website hacking [30], code vulnerability exploitation [31], debugging [32], and penetration testing [33]. In this paper, we focus on developing a standardized toolkits for evaluating LLM agents.
LLM Agent Benchmark.
@@
To evaluate the capabilities of LLM agents, several benchmarks have been developed. AgentBench [11] assesses LLMs across five diverse environments, including operating systems and databases, to evaluate reasoning and decision-making abilities. MLAgentBench [12] focuses on machine learning experimentation tasks, testing agents on tasks ranging from improving model performance to addressing research problems. SWE-bench [13] evaluates LLMs on real-world software issues sourced from GitHub, requiring models to generate patches that resolve described problems. SmartPlay [14] introduces a suite of games to test various capabilities of LLMs, such as planning and spatial reasoning. WebArena [15] provides a realistic web environment for building autonomous agents, enabling the assessment of LLMs in web-based tasks.
Cybersecurity-Specific Benchmarks.
@@
In the cybersecurity domain, specialized benchmarks have been introduced. Cybench [16] offers a framework for evaluating LLM agents on 40 professional-level Capture The Flag (CTF) tasks, encompassing a range of difficulties and scenarios. CyberMetric [17] presents a benchmark dataset based on retrieval-augmented generation to assess LLMs’ cybersecurity knowledge. SecEval [34] provides over 2,000 multiple-choice questions across various cybersecurity domains to evaluate foundation models’ knowledge. CyberSecEval [18] focuses on code vulnerability detection and exploitation, offering a comprehensive suite for assessing LLMs in secure coding tasks. These benchmarks facilitate targeted evaluations of LLMs in cybersecurity contexts. The closest work to ours is CyberBench [35], a benchmark focusing on Natural Language Processing (NLP) tasks related to cybersecurity.
DefenderBench.
@@
We introduce DefenderBench, a toolkit designed to evaluate LLM agents in interactive cybersecurity environments. Unlike existing benchmarks mentioned above that focus on specific tasks or domains, DefenderBench encompasses a broad range of cybersecurity-related tasks, covering offense, defense, and knowledge understanding. By integrating insights from general agent benchmarks and adversarial evaluation frameworks, DefenderBench aims to provide a comprehensive assessment platform for LLMs in cybersecurity contexts.
@@ 3.1 Computer Network Intrusion Simulation
In order to protect computer networks against attacks, many organizations conduct red-team network intrustion to proactively detect and remediate vulnerabilites before attackers do. We leverage the network intrusion simulation tool CyberBattleSim (CBS) [36] to evaluate the ability of LLM agents to identify vulnerabilities in a network. CyberBattleSim is parameterized by a fixed topology and a set of node vulnerabilities that agents can exploit to move laterally within the network. The goal of the attacker is to take ownership of the network by exploiting vulnerabilities in the computer nodes. We convert CyberBattleSim into a text-based game [37] which describes the currently discovered network as some structured text (i.e., JSON) and provides textual feedback in response to the agent’s actions. There are three action types for an attacker to interact with the network:
@@ 3.2 Malicious Content Detection
Malicious-Text: for this task, we utilize the dataset processed by Alvarado [38]. [^2] https://huggingface.co/datasets/ealvaradob/phishing-dataset This dataset incorporates two data sources, namely email and text messages, for malicious content detection. The entire dataset contains 20,137 samples labeled as $\{malicious,legitimate\}$. We follow the split of Alvarado [38], using 80% of the data as the training set and 20% as the test set. To reduce the cost of performing LLMs on our benchmark, we further randomly select 500 samples from the test split as our official test set in the benchmark. Additionally, we select 10 samples per class as the few-shot sampling pool for in-context learning (ICL) [39]. The metric used is the macro-F1 score.
Malicious-Web: This task assesses the ability of LLM agents to discriminate phishing from benign web sites. We use the Phishing Websites Dataset [40] as preprocessed by Alvarado [38] for malicious website detection. We also discard 144 samples which contain less than 100 characters as they are mostly outliers (e.g. page failed to load). The resulting dataset (15,612 samples) includes 10,220 labeled as legitimate and 5,392 as malicious. We follow the same 80%-20% split as Alvarado [38] and further uniformly subsample 500 test samples as our test set and 10 training samples per class as the few-shot sampling pool. We report the macro-F1 score for this task.
@@ 3.3 Cyber Threat Intelligence (CTI)
MCQA: This task assesses the ability of an LLM agent to understand recent threat intelligence and apply it to challenging questions. A multiple-choice question answering task that uses the CTI-MCQA dataset introduced by Alam et al. [41]. This dataset originally contains 2,500 questions, each associated with a CTI-related webpage or document. After filtering out questions linked to inaccessible webpage or document, we obtained 2,338 samples. We then randomly downsample and split these into a test set (500 questions) and a few-shot sampling pool (20 samples). Each question has four options, with only one correct answer. The metric for this task is macro-F1.
@@ 3.4 Code Vulnerability Detection
Vulnerable-CG: This task assesses the ability of LLM agents to detect vulnerabilities in code. We use the code vulnerability detection dataset included in CodeXGLUE [42], which is split into training (21,854 samples), validation (2,732 samples), and test sets (2,732 samples). Each sample is a C language function annotated with the label ‘vulnerable’ or ‘non-vulnerable’. Our test samples are 500 randomly selected samples from their test set. We also provide 10 training samples per class as the few-shot sampling pool. The agent’s performance is reported using the macro-F1 score.
Vulnerable-DV: we also include the Devign [43] dataset for code vulnerability detection in our benchmark. Zhou et al. [43] released two projects, FFmpeg and Qemu, comprising a total of 27,318 samples. We randomly sample 500 samples for our test set. Similarly, we include 10 training samples per class as the few-shot sampling pool and report the macro-F1 score as the evaluation metric.
@@ 3.5 Code Vulnerability Fixing
CVEFix: we use the CVEFix dataset [44] for the vulnerability fixing task. The original dataset contains 12,107 vulnerability fixing commits across 4,249 open-source projects. The dataset includes the source code before and after the changes. We only extract commits with the following conditions: (a) single method modification; (b) the commit is associated to a single CVE (Common Vulnerabilities and Exposures); (c) the programming languages is either: C, C++, Go, Java, JavaScript, PHP, Python, or Rust. As a result, we obtained 240 samples. We use all the samples as the test set for our benchmark. For this task, we provide the method’s source code before the commit and ask the agent to generate a new method that fixes any vulnerability. We report the CodeBLEU score [45] between the generated method and the method after the commit.
@@ 4.1 Modules
The DefenderBench toolkit automatically downloads the required datasets from their respective sources, shuffles the samples randomly according to a fixed random seed, and splits them into a test set and a few-shot sample pool for in-context learning. Once preprocessed, the datasets are cached locally. For network intrusion simulation, we install CyberBattleSim [36] as a dependency.
Task Environment.
@@
For each task, we set up a task environment that provides task-specific instructions (shown in Table 1), defines the action space for the agent, loads the relevant datasets and constructs few-shot examples if few-shot in-context learning is being conducted (more on this in section 5.3). For the detection, MCQA, and code-fixing tasks, each episode involves presenting the agent with a test sample. Each episode can run for up to five steps. If the agent fails to respond with the expected format, a feedback message is provided and the agent can try again until the episode ends. For the network intrusion task, each episode begins with an initialized network and can run for up to 100 steps to compromise the entire network. The LLM agent interacts with the task environment by providing a text action and the environment provides an observation in return. The observation describes the result of the given action and indicates whether the task has been completed. Additionally, the environment maintains a history of the actions taken by the agent and the corresponding feedback. The history can be provided to the agent as part of its context.
Agent Interface.
@@
Our DefenderBench is equipped with an LLM agent interface that enables users to integrate both open- and closed-weight LLMs. Users can also seamlessly incorporate their own agentic system to perform the tasks.
Execution.
@@
To evaluate LLM agents on DefenderBench, users can install our toolkit as a Python library. Through a terminal command, users can run all tasks or specify a particular task by using its shorthand name. Additionally, users can choose which LLM to use for the baseline agent. We have also integrated the Weights and Biases library into DefenderBench, [^3] https://wandb.ai/ enabling users to track and visualize their results seamlessly.
Metrics.
@@
We report on each task using its original metric as described in Section 3. Inspired by previous evaluation benchmarks like GLUE [46], we define a global metric called DefenderBench score, which represents the unweighted average of all task-specific metrics. The DefenderBench score provides an overall indication of performance on cybersecurity tasks.
Baseline Agent.
@@
To evaluate the out-of-the-box capability of LLMs in solving cybersecurity tasks, we experiment with a baseline agent with minimal scaffolding in this paper. As illustrated in Figure 1, we begin by providing to the agent a task instruction that explains the task, specifies the response format, and defines the action space. Table 1 shows the task instructions. At each step, the agent is given the trajectory of its prior actions along with the corresponding observations from the environment. At each step, the agent is asked to produce an action in the required format, which is then sent to the task environment to obtain an action observation. Based on this observation, we determine whether the episode should be terminated. If the episode continues, the observation is added to the system prompt as part of the historical trajectory.
@@ 5.1 Backbone LLMs
In our experiments, we use a variety of LLMs as the backbone of our agent. These include (1) open-weight models (Llama 3.1 [21], Llama 3.2, Llama 3.3, and Phi-3.5 [22]), (2) proprietary models (GPT-3.5, GPT-4-turbo, GPT-4o, GPT-4o-mini, Claude-3.5-haiku, and Claude-3.5-sonnet, Claude-3.7-sonnet), and (3) proprietary reasoning models (o1, o1-mini, o3, o4-mini, GPT-4.1, GPT-4.1-mini, and GPT-4.1-nano, Claude-3.7-sonnet-think).